A 4534 kg truck traveling with a velocity of 11 m/s due north collides head-on with a 1197 kg car traveling with a velocity of 12 m/s due south. The two vehicles stick together after the collision.
(a) What is the momentum of the truck prior to the collision?
(b) What is the momentum of the car prior to the collision?
(c) What is the magnitude and direction of the total momentum of the two vehicles after they collide?
Use positive for north and negative for south to indicate directions in your answers.
Answer #1

a)

Momentum of the truck before collision

Ptruck=MtruckVtruck=4534*11=49874 Kg-m/s

b)

Momentum of car before collision

Pcar=McarVcar=1197*(-12) =-14364 Kg-m/s

c)

By Conservation of momentum

Pf=49874 -14364 =35510 Kg-m/s

Direction is towards north